# MiKASA: Multi-Key-Anchor & Scene-Aware Transformer for 3D Visual Grounding

### v1

@@ Abstract
3D visual grounding involves matching natural language descriptions with their corresponding objects in 3D spaces. Existing methods often face challenges with accuracy in object recognition and struggle in interpreting complex linguistic queries, particularly with descriptions that involve multiple anchors or are view-dependent. In response, we present the MiKASA (Multi-Key-Anchor Scene-Aware) Transformer. Our novel end-to-end trained model integrates a self-attention-based scene-aware object encoder and an original multi-key-anchor technique, enhancing object recognition accuracy and the understanding of spatial relationships. Furthermore, MiKASA improves the explainability of decision-making, facilitating error diagnosis. Our model achieves the highest overall accuracy in the Referit3D challenge for both the Sr3D and Nr3D datasets, particularly excelling by a large margin in categories that require viewpoint-dependent descriptions. The source code and additional resources for this project are available on GitHub:
https://github.com/birdy666/MiKASA-3DVG
@@ 4 Experiment
[…]
M3DRef[40]ICCV 23 - - - - - 49.4% 55.6% 43.4% 42.3% 52.9%
@@ 4.2.3 Baseline Comparison
7Table 1 provides a comparative analysis of MiKASA against existing models in the Sr3D and Nr3D challenges[2]. MiKASA leads in overall accuracy for both Sr3D and Nr3D, achieving 75.2% and 64.4% respectively. Specifically, it demonstrates exceptional performance in the view-dependent category, achieving 70.4% in Sr3D and 65.4% in Nr3D, and significantly surpasses previous works. This performance underlines its capability to handle complex scenarios requiring changes in viewpoint, proving the effectiveness of our multi-key-anchor and features fusion strategy.
@@ 4.3 Ablation Studies
Spatial encoder -
[…]
Text-Spatial fusion -
Overall acc 64.4% 45.0% 62.4% 62.6%
Table 2: Ablations of the spatial module, results highlighting the essentiality of each component.
@@
[…]
PointNet++ 63.8%
PointNet++ & GCN 65.5%
PointNet++ & Self-Attention Based 70.8%

### v4

@@ Abstract
3D visual grounding involves matching natural language descriptions with their corresponding objects in 3D spaces. Existing methods often face challenges with accuracy in object recognition and struggle in interpreting complex linguistic queries, particularly with descriptions that involve multiple anchors or are view-dependent. In response, we present the MiKASA (Multi-Key-Anchor Scene-Aware) Transformer. Our novel end-to-end trained model integrates a self-attention-based scene-aware object encoder and an original multi-key-anchor technique, enhancing object recognition accuracy and the understanding of spatial relationships. Furthermore, MiKASA improves the explainability of decision-making, facilitating error diagnosis. Our model achieves the highest overall accuracy in the Referit3D challenge for both the Sr3D and Nr3D datasets, particularly excelling by a large margin in categories that require viewpoint-dependent descriptions. The source code and additional resources for this project are available on GitHub: https://github.com/dfki-av/MiKASA-3DVG
@@ 4 Experiment
BUTD-DETR[16]ECCV 22 67.0% 68.6% 63.2% 53.0% 67.6% 54.6% 60.7% 48.4% 46.0% 58.0%
NS3D[13]CVPR 23 62.7% 64.0% 59.6% 62.0% 62.7% - - - - -
@@
Table 1: Comparative analysis of the accuracy of our end-to-end solution against others on the Sr3D and Nr3D Challenges, highlighting MiKASA’s enhancements. Notably, in the view-dependent category, our solution demonstrates exceptional performance improvements, underscoring MiKASA’s superior accuracy and effectiveness over previous methods.
@@ 4.2.2 Evaluation Metrics
[…]
w/o Spatial encoder $45.0\%\pm 0.4\%$ $53.4\%\pm 0.6\%$ $36.9\%\pm 0.6\%$ $43.4\%\pm 0.4\%$ $45.8\%\pm 0.6\%$
w/o Feedforward layer $62.4\%\pm 0.2\%$ $69.0\%\pm 0.2\%$ $56.1\%\pm 0.6\%$ $62.4\%\pm 0.2\%$ $62.4\%\pm 0.2\%$
w/o Text-Spatial fusion $62.6\%\pm 0.1\%$ $69.0\%\pm 0.3\%$ $56.4\%\pm 0.2\%$ $61.8\%\pm 0.6\%$ $62.9\%\pm 0.1\%$
Table 2: Ablations of the spatial module on Nr3D, results highlighting the essentiality of each component.
@@ 4.2.3 Baseline Comparison
Table 1 provides a comparative analysis of MiKASA against existing models in the Sr3D and Nr3D challenges[2]. MiKASA leads in overall accuracy for both Sr3D and Nr3D, achieving 75.2% and 64.4% respectively. Specifically, it demonstrates exceptional performance in the view-dependent category, achieving 70.4% in Sr3D and 65.4% in Nr3D, and significantly surpasses previous works. This performance underlines its capability to handle complex scenarios requiring changes in viewpoint, proving the effectiveness of our multi-key-anchor and features fusion strategy.
@@
[…]
PointNet++ 63.8%
PointNet++ & GCN 65.5%
PointNet++ & Self-Attention Based 70.8%